OpenStudy (anonymous):

Find the area of the ring formed by the concentric circles defined by (x - 3)2 + (y – 2)2 = 16 and (x – 3)2 + (y – 2)2 = 144.

OpenStudy (zarkon):

\[128\pi\]

OpenStudy (zarkon):

\[\pi R^2-\pi r^2=\pi(R^2-r^2)\]

OpenStudy (anonymous):

Thanks Zarkon, for the equation and answer ^-^

OpenStudy (zarkon):

np